OpenStudy (anonymous):

Please help with this equation!! sqrt(x^2-3x)=2 is the answer x=1 or x=4, or both?

OpenStudy (anonymous):

\[\sqrt{x^2-3x}=2\]

OpenStudy (anonymous):

@binks

jimthompson5910 (jim_thompson5910):

Let's check each one individually

OpenStudy (anonymous):

1 doesn't work right? it's only four, right?

jimthompson5910 (jim_thompson5910):

Checking x = 1: \[\large \sqrt{x^2-3x}=2\] \[\large \sqrt{(1)^2-3(1)}=2\] \[\large \sqrt{1-3(1)}=2\] \[\large \sqrt{1-3}=2\] \[\large \sqrt{-2}=2\] \[\large \sqrt{-1*2}=2\] \[\large \sqrt{-1}*\sqrt{2}=2\] \[\large i\sqrt{2}=2\] That's FALSE. So x = 1 is NOT a solution.

OpenStudy (anonymous):

I was trying to solve the equation instead of plugging in the values! I'm so dumb. You've been very helpful jim thompson.

jimthompson5910 (jim_thompson5910):

Checking x = 4: \[\large \sqrt{x^2-3x}=2\] \[\large \sqrt{(4)^2-3(4)}=2\] \[\large \sqrt{16-3(4)}=2\] \[\large \sqrt{16-12}=2\] \[\large \sqrt{4}=2\] \[\large 2=2\] That's TRUE. So x = 4 is defintinitely a solution.

jimthompson5910 (jim_thompson5910):

Well you should have solved first to get list of potential solutions

jimthompson5910 (jim_thompson5910):

After you solve, you test each potential solution to see if it is a true solution or not.

OpenStudy (anonymous):

You're right! I will do that from now on